manny a single taxpayer erans 66400 per year in taxable income and an additional 12140 per year in city of boston bonds.

if manny earns an additional 36400 in taxable income 2020 what is his marginal tax rate on this income?
Answer #1
Tax on $66,400 $10,398 =4617.5+((66400-40125)*22%)
Additional Income 36,400
Revised Taxable Income        102,800 =66400+36400
Tax on $ 102,800 $18,751.50 =14605.5+((102800-85525)*24%)
Marginal tax rate on this income 22.95% =(18751.5-10398)/36400
or
23%